OpenStudy (anonymous):

********HELP PLEASE************** A local eye clinic has two doctors. Dr. Kutallot recommends glass 60% of the time and corrective surgery 40% percent of the time. Dr. Lenzie recommends glasses 70% of the time and 30% of the time. 60% of the patients are treated by Dr. Lenzie. No patients are treated by both doctors. If a patient is selected random, what is the probability that surgery is recommended? 18% 34% 35% 58% 70%

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

how did you get 0.4 and 0.3

OpenStudy (anonymous):

got it .. answer is 34%.

OpenStudy (anonymous):

\[(40×0.4)+(60×0.3)=34\]

OpenStudy (anonymous):

how did you get the 0.4 and 0.3

OpenStudy (anonymous):

can you help with one more please

OpenStudy (anonymous):

if performances on AP test are independent and the probability of passing an AP is 0.3 then the probability of passing three AP test is 0.027 0.090 0.300 0.900 0.333

OpenStudy (anonymous):

lets say there are 100 patients. out of which 40 patients go to Dr. kuttalot. and he recommends corrective surgery to 40% patients (16 patients). 60 patients meet Dr. Lanze and out of he/she recommends corrective surgery to 30% (18 patients). hence total patients recommended with corrective surgery out of 100 total patients is = 16 +18 = 34

OpenStudy (anonymous):

\[0.3\times0.3\times0.3 = 0.027\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

didnt award me medals :/

OpenStudy (anonymous):

how do you do the medal

OpenStudy (anonymous):

well thanks! I got the medal.